OpenStudy (anonymous):

Verify the following. 1) Solution x^2+y^2=Cy Differential Equation y'=(2xy)/(x^2-y^2)

myininaya (myininaya):

x^2+y^2=Cy ---------------------------------------- (x^2+y^2=Cy)' 2x+2yy'=Cy' solve for y' 2x=Cy'-2yy' 2x=y'(C-2y) 2x/(C-2y)=y' y'=2x/(C-2y)

myininaya (myininaya):

but C=(x^2+y^2)/y

myininaya (myininaya):

\[y'=\frac{2x}{\frac{x^2+y^2}{y}-2y}\]

myininaya (myininaya):

\[y'=\frac{2x}{\frac{x^2+y^2}{y}-2y}*\frac{y}{y}\]

myininaya (myininaya):

\[y'=\frac{2xy}{x^2+y^2-2y^2}\]

myininaya (myininaya):

\[y'=\frac{2xy}{x^2-y^2}\]

OpenStudy (anonymous):

you always provide the most elaborate explanations. Thanks, myininaya :]

myininaya (myininaya):

you're so sweet

OpenStudy (anonymous):

wow!